OpenStudy (anonymous):

Verify. cscx-tan(x/2)=cotx

OpenStudy (anonymous):

I want to do LHS--->RHS

OpenStudy (misty1212):

HI!!

OpenStudy (amorfide):

do you know your identities?

OpenStudy (misty1212):

i guess you have to dredge up the formula for \[\tan(\frac{x}{2})\] which i can never remember

OpenStudy (misty1212):

actually now that i look closer, i am pretty sure it is NOT and identity lets try it with \(\frac{\pi}{3}\)

OpenStudy (anonymous):

cscx = 1/sinx and tan(x/2) = 1-cosx/2 = sinx/1+cosx

OpenStudy (misty1212):

\[\csc(\frac{\pi}{3})=2\]\[\tan(\frac{\pi}{6})=\frac{1}{\sqrt3}\]and \[2-\frac{1}{\sqrt3}\neq \cot(\frac{\pi}{3})=\frac{1}{\sqrt3}\] so you cannot prove it

OpenStudy (anonymous):

Thank you.